OpenStudy (anonymous):

From a boat on the lake, the angle of elevation to the top of a cliff is 12°50'. If the base of the cliff is 1366 feet from the boat, how high is the cliff (to the nearest foot)?

OpenStudy (amistre64):

1366(tan(12'50'')) '=degree and ''=minute for this one :)

OpenStudy (anonymous):

okay :)

OpenStudy (amistre64):

(79/6) degrees ??

OpenStudy (amistre64):

319.554 is what the calculator spits out :) that is if 79/6 is a good number. 12 and 50/60 = 12 5/6 = (72+5)/6 = 77/6 gotta recalc it :)

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

311.183 lets try that one on fer size :)

OpenStudy (anonymous):

i got 311.09

OpenStudy (amistre64):

close enough....311 is prolly the answer then

OpenStudy (anonymous):

:) thank you can you help me with my next one

OpenStudy (amistre64):

maybe :) is it an easy one?

OpenStudy (anonymous):

its kind of the same as this

OpenStudy (amistre64):

lets here it

OpenStudy (anonymous):

From a boat on the river below a dam, the angle of elevation to the top of the dam is 30°25'. If the dam is 1065 feet above the level of the river, how far is the boat from the base of the dam (to the nearest foot)?

OpenStudy (amistre64):

these are basically tan problems

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

[30(60)+5]/60 is going to be the number for my degrees.. tan(n) = 1065/x x = 1065/tan(n)...

OpenStudy (amistre64):

1838.453 double check that...

OpenStudy (amistre64):

1838.453 its what I get :)

OpenStudy (anonymous):

1813 ft?

OpenStudy (amistre64):

nope... messed up someplace... my number for degrees specifically..

OpenStudy (amistre64):

[30(60)+ 25] /60

OpenStudy (amistre64):

recalculating....

OpenStudy (amistre64):

1814.03 is what i get with the right numbers :)

OpenStudy (anonymous):

i did it that way and got 1838.7

OpenStudy (anonymous):

wow lol

OpenStudy (amistre64):

I added a "5" it aint 5/60 minutes... its 25/60 minutes so I had to get the right number for degrees :)

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

im sure its 1814, or close to that :)

OpenStudy (anonymous):

thank you :)